OpenStudy (anonymous):

In a scale drawing of Samantha's apartment, the ratio is 1 in = 2 ft. The length of the living room on the drawing is 10 in and the width is 4 in. What is the perimeter of Samantha's living room?

OpenStudy (anonymous):

a 56 ft b 40 ft c 28 ft

OpenStudy (anonymous):

so lenght is 10 * 2 = 20ft and wiidth is 4 * 2 = 8ft Permeter = 2*(length + width)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

\[P = 2(20+8) = 56ft\]